A 2.00 kg mass is placed on the top of a hill 10.0 m high. Assume no energy is being lost through heat or friction. If the spring compresses 0.500 m when the ball comes to a stop, what is the spring constant?


1
Expert's answer
2021-05-14T18:22:33-0400

"mgh=kx^2/2\to k=2mgh/x^2=2\cdot2\cdot9.81\cdot 10/0.5^2=1570\ (N/m)" . Answer